# The Renormalization Group for Large-Scale Structure: Origin of Galaxy Stochasticity

### v1

@@ Abstract
The renormalization group equations for large-scale structure (RG-LSS) describe how the bias and stochastic (noise) parameters—both of matter and biased tracers such as galaxies—evolve as a function of the cutoff $\Lambda$ of the effective field theory. In previous work, we derived the RG-LSS equations for the bias parameters using the Wilson-Polchinski framework. Here, we extend these results to include stochastic contributions, corresponding to terms in the effective action that are higher order in the current $J$. We show that the RG equations exhibit an interesting, previously unnoticed structure at all orders in $J$, which implies that a single nonlinear bias term immediately generates all stochastic moments through RG evolution. We then derive the nonlinear RG evolution of the (leading-derivative) stochastic parameters for all $n$-point functions, and show that this evolution is controlled by a different, lower scale than the nonlinear scale. This has implications for the optimal choice of the renormalization scale when comparing the theory with data to obtain cosmological constraints.
@@ 1.1 Summary of main results
Note that the set of operators $O$ appearing here is always the same, and terms with $m=1$ are nothing but the bias expansion with $C_{O}^{(1)}\equiv b_{O}$. We explain how $S_{\rm eff}$ generates the $n$-point functions of galaxy clustering after Eq. (2.7). Expanding the operator functionals $O[\delta^{(1)}_{\Lambda}]$ in Eq. (1.2) [as discussed in more detail below, cf. Eq. (1.18)], we recover the effective action introduced by [3]
@@
with each $\mathcal{K}^{(m)j_{1}\ldots j_{m}}_{i_{1}\ldots i_{n}}(\Lambda)$ representing a sum over operator kernels with associated coefficients $C_{O}^{(m)}$. [^3] Specifically, we have $\mathcal{K}^{(m)j_{1}\ldots j_{m}}_{i_{1}\ldots i_{n}}(\Lambda)=(2\pi)^{3}\delta_{\rm D}(\bm{k}_{j_{1}\ldots j_{m}}+\bm{k}_{i_{1}\ldots i_{n}})\sum_{O}C^{(m)}_{O}(\Lambda)K^{(n)}_{O}(\bm{k}_{i_{1}},\ldots,\bm{k}_{i_{n}}),$ (1.4) where the sum runs over all bias operators contributing at order $n$ and $K^{(n)}_{O}$ is defined in Eq. (1.18). The fact that Ref. [3] considered a special case of biased tracer, matter, is not important for the structure of the equations. Ref. [3] then derive RG equations for the coefficient tensors $\mathcal{K}^{(m)}(\Lambda)$ [Eq. (4.22) there; notice that in their notation, $n$ and $m$ are swapped and their $K$ corresponds to our $\mathcal{K}$]. For clarity, let us drop all indices and combinatorial factors in the following. The structure of the RGE derived by Ref. [3] is
@@
We will argue in this paper that the second term in Eq. (1.5) in fact vanishes in general due to kinematics. Instead, the general RG-LSS equation becomes
@@
where the series continues to all orders in $dP_{\rm L}^{\Lambda}/d\Lambda$. While one might think that terms with higher powers of $dP_{\rm L}^{\Lambda}/d\Lambda$ are suppressed since one is considering a narrow shell in momentum space, this is not the case as we will show (hint: all terms in Eq. (1.6) are one-loop diagrams). Note that for $m=1$ (bias terms), only the first term in both Eq. (1.5) and Eq. (1.6) contributes, and we recover the RGE for bias parameters studied in detail in Paper 1.
Clearly, the structure of Eq. (1.6) is quite different from, and more complex than that of Eq. (1.5). We will explain in detail how Eq. (1.6) arises, and what the physical ramifications are. The main results are the following:
@@
  • Finally, we express Eq. (1.6) as an explicit set of nonlinear ordinary differential equations for the stochastic amplitudes, given the bias coefficients $b_{O}(\Lambda)$ as a function of scale. This nonlinearity leads to interesting behavior under RG flow, and ramifications for the choice of the renormalization scale in EFT analyses, which we investigate in Sec. 4.
@@ 1.2 Notation
We stress again that these relations mainly serve to give meaning to the coefficients $P_{\epsilon,{\mathbb{1}}},P_{\epsilon,O}$ and $C_{\epsilon,O}^{(m)}$ in general; the partition function is completely written in terms of these coefficients and without reference to $\epsilon(\bm{k})$.
@@ 2 The general EFT partition function for bias and stochastic parameters
beyond the leading stochastic contributions considered there. Here, the sum runs over all bias operators $O$ that are relevant at a given order in perturbations and derivatives. [^4] For other works on the large-scale structure partition function, see also [19, 20, 21, 22, 13]. Notice that Eq. (2.1) includes a nonlinear coupling between the field and the current $J$, via the operators $O[\delta^{(1)}_{\Lambda}]$. Terms that couple nonlinearly to the current are usually referred to as “composite operators” (see [23, 24] and the books [25, 26]). As discussed in [23], the renormalization of those composite operators will lead to terms that are higher-order in $J$. As we will see in Sec. 3, those higher-order stochastic contributions are necessary to obtain a partition function that remains consistent under the RG flow. Thus, the general effective action, writing explicitly the $\mathcal{O}[J_{\Lambda}^{2}\delta^{(1)}_{\Lambda},\ J_{\Lambda}^{3}]$ terms of Eq. (2.1), is given by a sum over local-in-space products of currents and bias operators
@@
where we include the trivial constant operator $O(\bm{x})={\mathbb{1}}$ in the set of bias operators, which is more convenient when going to higher order in $J$. We can then identify $b_{O}(\Lambda)\equiv C^{(1)}_{O}(\Lambda)$ as the bias parameters [^5] Notice that the tadpole coefficient $b_{\mathbb{1}}=C^{(1)}_{{\mathbb{1}}}=0$ after setting $\langle\delta_{g}\rangle=0$. at the scale $\Lambda$, and $P_{\epsilon,O}(\Lambda)\equiv C^{(2)}_{O}(\Lambda)$ as the second-order “Gaussian” stochastic amplitudes, with $P_{\epsilon,{\mathbb{1}}}$ corresponding to the well-known stochastic or “shot noise” contribution to the galaxy power spectrum. Similarly, $P_{\epsilon,\delta}$ quantifies the leading coupling of stochasticity to the density (“density-dependent shot noise”). However, Eq. (2.2) also includes non-Gaussian stochastic contributions, starting at $m=3$ with $B_{\epsilon,{\mathbb{1}}}\equiv C^{(3)}_{{\mathbb{1}}}$ as the purely stochastic contribution to the galaxy bispectrum.
Notice that the interactions among the $J$ and $O$ in Eq. (2.2) are purely local. This is a consequence of working at leading order in derivatives for the stochastic contributions. Higher-derivative contributions, such as $J^{m-1}\nabla^{2}J$ and $J^{m-2}\partial_{i}J\partial^{i}J$ are also present in general, and are expected to be controlled by the same scale as that controlling the higher-derivative bias operators. We will neglect them in the following, however.
@@ 3.1 Integrating out a momentum shell via Wilson-Polchinski
is a symmetry factor in which $\mathcal{N}(a,\{i_{1},\dots,i_{p}\})$ is the number of times $a$ appears in $\{i_{1},\dots,i_{p}\}$; hence, $g^{1}_{1}=1$, $g^{11}_{2}=1$, $g^{2}_{2}=1$, $g^{111}_{3}=1$, $g^{12}_{3}=g^{21}_{3}=3$, $g^{3}_{3}=1$, and so forth. Notice that we isolated a factor $\frac{1}{m!}\prod_{a=1}^{m}\left[\frac{1}{\mathcal{N}(a,\{i_{1},\dots,i_{p}\})!}\right]$ in the first line of Eq. (3.15) for reasons that will become clear in Sec. 3.4. Eq. (3.15) is the master equation that will guide the following sections. One can trivially recover Eq. (3.7), Eq. (3.8) and Eq. (3.9) for $m=1,2,3$ respectively. On the other hand, it is important to stress the difference with respect to Eq. (4.22) of [3], which only includes the contributions of the type $S^{2}_{O}$ and $S^{11}_{OO^{\prime}}$. Crucially, as we will see later and already anticipated in Eq. (3.15), we find that the only relevant contributions come from shell contributions $\mathcal{S}_{O_{1}\dots O_{p}}^{2\dots 2}$, while the $S^{11}_{OO^{\prime}}$ term is actually suppressed due to its kinetic structure and the orthogonality with the current $J_{\Lambda}$.
@@
We dedicate App. A.2 to discuss how the contributions coming from a single operator but expanded in higher orders of shell fields ($\mathcal{S}_{O^{\prime}}^{4}$, $\mathcal{S}_{O^{\prime}}^{6}$, $\dots$) are suppressed by extra factors of $\lambda$ compared to $\mathcal{S}_{O^{\prime}}^{2}$. Since we can take the limit $\lambda\to 0$ without loss of generality of the RG flow, we can therefore drop these terms, and it is sufficient to consider $\mathcal{S}_{O^{\prime}}^{2}$ in $\zeta^{(m)}[J_{\Lambda},\delta^{(1)}_{\Lambda}]$.
@@
Instead, the leading contribution comes from the $\mathcal{S}_{O^{\prime}O^{\prime\prime}}^{22}$ term. This was in fact already anticipated in Eq. (3.13) of [13]. As is apparent from the the top right diagram in Fig. 2, the kinematic suppression does not apply in this case, as two modes with momenta in the shell can couple to produce an external momentum $k\ll\Lambda$. In fact, this is a specific case of the generic conclusion of App. B.2 that the terms of the type $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}^{22\dots 2}$ are the leading sources in general. To exemplify this, we show the corresponding tree-level and 1-loop diagrams for the $\mathcal{S}_{OO^{\prime}O^{\prime\prime}}$ terms in Fig. 3.
@@
Notice the specific scaling of each contribution with different powers of $\left[P_{\rm L}(\Lambda)\right]^{p-1}$, which is required by dimensional analysis [Eq. (2.5)].
@@ 3.5 Summary
  • If all source terms to the RG equations were of the $\mathcal{S}^{2}_{O}$ type, there would be no coupling between terms with different powers in the current, $J^{m}$. In particular, no stochasticity ($m\geq 2$) would be generated by nonlinear bias ($m=1$). However, we find that terms of the form $\mathcal{S}_{O_{1}\dots O_{m}}^{2\dots 2}$ generate contributions of similar relevance as $\mathcal{S}^{2}_{O}$ to all higher $m$. That is, a single nonlinear bias term $b_{\delta^{2}}(\Lambda_{\ast})$ immediately generates all stochastic moments when running to a scale $\Lambda<\Lambda_{\ast}$. This finding differs substantially from the conclusions of [3], who only considered $\mathcal{S}_{O}^{2}$ and $\mathcal{S}_{O_{1}O_{2}}^{11}$ as relevant contributions. We find $\mathcal{S}_{O_{1}O_{2}}^{11}$ to be kinematically suppressed, while $\mathcal{S}_{O_{1}O_{2}}^{22}$, and its generalization to $2\to m$ operators, is unsuppressed. [^8] Another elucidating and perhaps more familiar example of this type of contribution is the renormalization of $\delta^{(2)}\delta^{(2)}$ that appears for the matter field, which sources a higher-derivative stochastic operator proportional to $k^{4}$, and therefore absorbed by a term $\left[\nabla^{2}J(\bm{x})\right]^{2}$ in the effective action.

### v2

@@ Abstract
The renormalization group equations for large-scale structure (RG-LSS) describe how the bias and stochastic (noise) parameters—both of matter and biased tracers such as galaxies—evolve as a function of the cutoff $\Lambda$ of the effective field theory. In previous work, we derived the RG-LSS equations for the bias parameters using the Wilson-Polchinski framework. Here, we extend these results to include stochastic contributions, corresponding to terms in the effective action that are higher order in the current $J$. We derive the general local interaction terms that describe stochasticity at all orders in perturbations, and a closed set of nonlinear RG equations for their coefficients. These imply that a single nonlinear bias term generates all stochastic moments through RG evolution. Further, the evolution is controlled by a different, lower scale than the nonlinear scale. This has implications for the optimal choice of the renormalization scale when comparing the theory with data to obtain cosmological constraints.
@@ 1 Introduction
We provide the most general RG running of all leading-in-derivative EFT parameters [see Eq. (3.36)], elucidating in detail how bias coefficients source stochastic terms. This result could be used in practice as an extra validation (both at field level and using $n$-point correlation functions) for the running of the EFT parameters. In addition, our framework can straightforwardly renormalize any higher-order $n$-point functions (including their respective stochastic contributions). Moreover, our RG approach keeps the smoothing scale $\Lambda$ explicit, going beyond the usual $n$-point function renormalization scheme considered in other works [8] and proven in Sec. 3 of Paper 1 to correspond the limit $\Lambda\to 0$ of the RG equations.
@@ 1.1 Summary of main results
Note that the set of operators $O$ appearing here is always the same, and terms with $m=1$ are nothing but the bias expansion with $C_{O}^{(1)}\equiv b_{O}$. We explain how $S_{\rm eff}$ generates the $n$-point functions of galaxy clustering after Eq. (2.7). Expanding the operator functionals $O[\delta^{(1)}_{\Lambda}]$ in Eq. (1.2) [as discussed in more detail below, cf. Eq. (1.19)], we recover the same structure as the effective action introduced by [3]
@@
with each $\mathcal{K}^{(m)j_{1}\ldots j_{m}}_{i_{1}\ldots i_{n}}(\Lambda)$ representing a sum over operator kernels with associated coefficients $C_{O}^{(m)}$. [^3] Specifically, we have $\mathcal{K}^{(m)j_{1}\ldots j_{m}}_{i_{1}\ldots i_{n}}(\Lambda)=(2\pi)^{3}\delta_{\rm D}(\bm{k}_{j_{1}\ldots j_{m}}+\bm{k}_{i_{1}\ldots i_{n}})\sum_{O}C^{(m)}_{O}(\Lambda)K^{(n)}_{O}(\bm{k}_{i_{1}},\ldots,\bm{k}_{i_{n}}),$ (1.4) where the sum runs over all bias operators contributing at order $n$ and $K^{(n)}_{O}$ is defined in Eq. (1.19). The fact that Ref. [3] considered a special case of biased tracer, matter, is not important for the structure of the equations. Ref. [3] derive RG equations for the coefficient tensors $\mathcal{K}^{(m)}(\Lambda)$ [Eq. (4.22) there; notice that in their notation, $n$ and $m$ are swapped and their $K$ corresponds to our $\mathcal{K}$]. For clarity, let us drop all indices and combinatorial factors in the following. The structure of the RGE derived by Ref. [3] is
@@
The two contributions in Eq. (1.5) are precisely the two terms identified in the original paper by Polchinski [12], which each involve precisely one propagator in the momentum shell. The downside that comes with this simple structure is that one obtains nonlocal interactions. While these can be directly expanded in derivatives when considering the first “loop” term in Eq. (1.5), the second “tree-level” cannot be directly expanded. In the simplest nontrivial case for our effective action, this interaction looks like [Eq. (B.7)]
$\int_{\bm{x}}J_{\Lambda}(\bm{x})\delta^{(1)}_{\Lambda}(\bm{x})\int_{\bm{x}^{\prime}}J_{\Lambda}(\bm{x}^{\prime})\delta^{(1)}_{\Lambda}(\bm{x}^{\prime})\xi_{\rm shell}(\bm{x}-\bm{x}^{\prime}),$ (1.6)
where $\xi_{\rm shell}(r)$ is the correlation function of modes in the momentum shell (i.e. the Fourier transform of the shell propagator). This type of nonlocal term can be shown to vanish in the limit that the external momenta contributing to the current $J_{\Lambda}$ are much less than the scale $\Lambda$ (App. B.1); however, in the Polchinski approach we cannot directly set the external momenta to be small, because they include the momentum shell that is to be integrated out in the next step.
The question then is, whether and if so, how, the local interactions in Eq. (1.2) arise. We show in this paper that it is possible to work with soft external momenta in the RG flow from the beginning, resulting in local interactions from the outset. The price to pay for this significant simplification is that one now needs to consider contributions with higher powers of shell propagators in the RG flow. Nevertheless, we succeed in arriving at a closed set of (nonlinear) RGE, the general local RG-LSS equations in the form of
@@
where the series continues to all orders in $dP_{\rm L}^{\Lambda}/d\Lambda$. While one might think that terms with higher powers of $dP_{\rm L}^{\Lambda}/d\Lambda$ are suppressed since one is considering a narrow shell in momentum space, this is not the case as we will show; conversely, all other possible types of diagrams are suppressed (the reason: the terms in Eq. (1.7) are all one-loop diagrams). Note that for $m=1$ (bias terms), only the first term in both Eq. (1.5) and Eq. (1.7) contributes, and we recover the RGE for bias parameters studied in detail in Paper 1.
Clearly, the structure of Eq. (1.7) is quite different from, and more complex than that of Eq. (1.5). However, unlike Eq. (1.5), we can work with local interactions from the beginning. We will explain in detail how Eq. (1.7) arises, and what the physical ramifications are. The main results are the following:
@@
  • Finally, we express Eq. (1.7) as an explicit set of nonlinear ordinary differential equations for the stochastic amplitudes and bias parameters $C_{O}^{(m)}(\Lambda)$. This nonlinearity leads to interesting behavior under RG flow, and ramifications for the choice of the renormalization scale in EFT analyses, which we investigate in Sec. 4.
It appears much more difficult to arrive at such explicit results when dealing with the Polchinski RG Eq. (1.5) with nonlocal interactions instead.
In short, we present the most general RG running of all EFT parameters [see Eq. (3.36)], explaining how stochastic terms are sourced via the running of the EFT smoothing scale and showing that the effective action Eq. (1.2) is closed under renormalization.
@@ 1.2 Notation
We stress again that these relations mainly serve to give meaning to the coefficients $P_{\epsilon,{\mathbb{1}}},P_{\epsilon,O}$ and $C_{\epsilon,O}^{(m)}$ in general; the partition function is completely written in terms of these coefficients and without reference to $\epsilon(\bm{k})$. An important result we obtain is that the action Eq. (2.2), including stochasticity, is closed under renormalization. Therefore the treatment of stochasticity that we provide in this work covers its most general properties, regardless of the LSS tracer considered, and also directly determines the stochastic contributions to all $n$-point functions.
@@ 2 The general EFT partition function for bias and stochastic parameters
beyond the leading stochastic contributions considered there. Here, the sum runs over all bias operators $O$ that are relevant at a given order in perturbations and derivatives. [^4] For other works on the large-scale structure partition function, see also [19, 20, 21, 22, 13]. Notice that Eq. (2.1) includes a nonlinear coupling between the field $\delta^{(1)}_{\Lambda}$ and the current $J$, via the operators $O[\delta^{(1)}_{\Lambda}]$. Terms that couple nonlinearly to the current are usually referred to as “composite operators” (see [23, 24] and the books [25, 26]). As discussed in [23], the renormalization of those composite operators will lead to terms that are higher-order in $J$. As we will see in Sec. 3, those higher-order stochastic contributions are necessary to obtain a partition function that remains consistent under the RG flow. Thus, the general effective action, writing explicitly the $\mathcal{O}[J_{\Lambda}^{2}\delta^{(1)}_{\Lambda},\ J_{\Lambda}^{3}]$ terms of Eq. (2.1), is given by a sum over local-in-space products of currents and bias operators
@@
where we include the constant operator $O(\bm{x})={\mathbb{1}}$ in the set of bias operators, which is more convenient when going to higher order in $J$. We can then identify $b_{O}(\Lambda)\equiv C^{(1)}_{O}(\Lambda)$ as the bias parameters [^5] Notice that the tadpole coefficient $b_{\mathbb{1}}=C^{(1)}_{{\mathbb{1}}}=0$ after setting $\langle\delta_{g}\rangle=0$. at the scale $\Lambda$, and $P_{\epsilon,O}(\Lambda)\equiv C^{(2)}_{O}(\Lambda)$ as the second-order “Gaussian” stochastic amplitudes, with $P_{\epsilon,{\mathbb{1}}}$ corresponding to the well-known stochastic or “shot noise” contribution to the galaxy power spectrum. Similarly, $P_{\epsilon,\delta}$ quantifies the leading coupling of stochasticity to the density (“density-dependent shot noise”). However, Eq. (2.2) also includes non-Gaussian stochastic contributions, starting at $m=3$ with $B_{\epsilon,{\mathbb{1}}}\equiv C^{(3)}_{{\mathbb{1}}}$ as the purely stochastic contribution to the galaxy bispectrum.
Notice that the interactions among the $J$ and $O$ in Eq. (2.2) are purely local. As noted in Sec. 1.1, the fact that we can write a local expression for the partition function that is closed under RG flow is a nontrivial result. Exact locality in fact is only obtained when working at leading order in derivatives for the stochastic contributions. Higher-derivative contributions, such as $J^{m-1}\nabla^{2}J$ and $J^{m-2}\partial_{i}J\partial^{i}J$ are also present in general, and are expected to be controlled by the same scale as that controlling the higher-derivative bias operators. We will neglect them in the following, however.
@@ 3.1 Integrating out a momentum shell via Wilson-Polchinski
is a symmetry factor in which $\mathcal{N}(a,\{i_{1},\dots,i_{p}\})$ is the number of times $a$ appears in $\{i_{1},\dots,i_{p}\}$; hence, $g^{1}_{1}=1$, $g^{11}_{2}=1$, $g^{2}_{2}=1$, $g^{111}_{3}=1$, $g^{12}_{3}=g^{21}_{3}=3$, $g^{3}_{3}=1$, and so forth. Notice that we isolated a factor $\frac{1}{m!}\prod_{a=1}^{m}\left[\frac{1}{\mathcal{N}(a,\{i_{1},\dots,i_{p}\})!}\right]$ in the first line of Eq. (3.15) for reasons that will become clear in Sec. 3.4. Eq. (3.15) is the master equation that will guide the following sections. One can trivially recover Eq. (3.7), Eq. (3.8) and Eq. (3.9) for $m=1,2,3$ respectively. On the other hand, it is important to stress the difference with respect to Eq. (4.22) of [3], which only includes the contributions of the type $S^{2}_{O}$ and $S^{11}_{OO^{\prime}}$. As discussed in Sec. 1.1, our goal here is to take the external momenta in the current to zero, to obtain a partition function that only has local interactions. In that case, as we will see later and already anticipated in Eq. (3.15), we need to include diagrams with more than one shell propagator, but only the specific contributions $\mathcal{S}_{O_{1}\dots O_{p}}^{2\dots 2}$. In the same limit of low external momenta, the $S^{11}_{OO^{\prime}}$ term is instead suppressed due to kinematics and the orthogonality with the current $J_{\Lambda}$.
@@
We dedicate App. A.2 to discuss how the contributions coming from a single operator but expanded in higher orders of shell fields (2- and higher-loop terms: $\mathcal{S}_{O^{\prime}}^{4}$, $\mathcal{S}_{O^{\prime}}^{6}$, $\dots$) are suppressed by extra factors of $\lambda$ compared to $\mathcal{S}_{O^{\prime}}^{2}$. Since we can take the limit $\lambda\to 0$ without loss of generality of the RG flow, we can therefore drop these terms, and it is sufficient to consider $\mathcal{S}_{O^{\prime}}^{2}$ in $\zeta^{(m)}[J_{\Lambda},\delta^{(1)}_{\Lambda}]$.
@@
In the limit of soft external moments, the leading contribution instead comes from the $\mathcal{S}_{O^{\prime}O^{\prime\prime}}^{22}$ term. This was in fact already anticipated in Eq. (3.13) of [13]. As is apparent from the the top right diagram in Fig. 2, the kinematic suppression does not apply in this case, as two modes with momenta in the shell can couple to produce an external momentum $k\ll\Lambda$. In fact, this is a specific case of the generic conclusion of App. B.2 that the terms of the type $\mathcal{S}_{O_{1}O_{2}\dots O_{p}}^{22\dots 2}$ are the leading sources in general. To exemplify this, we show the corresponding tree-level and 1-loop diagrams for the $\mathcal{S}_{OO^{\prime}O^{\prime\prime}}$ terms in Fig. 3.
@@
Notice the specific scaling of each contribution with different powers of $\left[P_{\rm L}(\Lambda)\right]^{p-1}$, which is required by dimensional analysis [Eq. (2.5)]. We emphasize once more the importance of Eq. (3.36): it governs the general RG flow of all leading-in-derivatives EFT parameters. It makes manifest how any coefficient (including bias and stochastic parameters) of the EFT of LSS is sourced by a variation in the smoothing scale $\Lambda$. Whether the RG running can lead to more (non-perturbative) information, such as in the context of quantum field theory, will be explored in a further publication. We refer to Sec. 4 of Paper 1 for a comparison to other renormalization schemes.
@@ 3.5 Summary
  • If all source terms to the RG equations were of the $\mathcal{S}^{2}_{O}$ type, there would be no coupling between terms with different powers in the current, $J^{m}$. In particular, no stochasticity ($m\geq 2$) would be generated by nonlinear bias ($m=1$). However, we find that terms of the form $\mathcal{S}_{O_{1}\dots O_{m}}^{2\dots 2}$ generate contributions of similar relevance as $\mathcal{S}^{2}_{O}$ to all higher $m$. Thus, according to our RGE system, a single nonlinear bias term $b_{\delta^{2}}(\Lambda_{\ast})$ immediately generates all stochastic moments when running to a scale $\Lambda<\Lambda_{\ast}$. In contrast, Ref. [3] found that higher-order stochastic terms are generated by successive “rounds” of the RG flow, as they only considered $\mathcal{S}_{O}^{2}$ and $\mathcal{S}_{O_{1}O_{2}}^{11}$ contributions. As discussed in Sec. 1.1, the difference arises because the intermediate state in the approach of [3] involves nonlocal interactions, while our RGE are written directly in terms of the local interaction terms that emerge in the IR limit.
In short, we have derived the most general running of the EFT of LSS coefficients in Eq. (3.36). This result can be straightforwardly used (both at field-level and by $n$-point functions) to validate the running of the freee EFT coefficients. It can also be used to renormalize any $n$-point function at any loop order including stochastic terms. Our result goes beyond the usual $n$-point function renormalization considered by other works considering the full RG running of the parameters, with the $n$-point function normalization being the $\Lambda\to 0$ limit of the RG running (see Sections 3 and 4 of [10] for a connection between both schemes).